The area of a rectangular wall of a barn is 108 square feet. Its length is 6 feet longer than twice its width. Find the length and width of the wall of the barn?
Answer by oberobic(2304)

L * W = 108 sq ft
Length is 6 ft greater than twice the width.
L = 2W + 6
Substituting
(2W+6)*W = 108
2W^2 + 6W = 108
Divide both sides by 2
W^2 + 3W = 54
W^2 + 3W - 54 = 0
Factor
(W +9)(W -6) = 0
Roots are W=-9 and W=6.
A negative width is impossible, so W=6
L = 2W + 6
L = 2(6) + 6
L = 18
.
Always check your answer.
L*W = ??
18 * 6 = 108
Correct.
.
Done